ghostmasterTy:

Simplify these improper fractions. Remember to divide the denominator into the numerator to simplify. _____ 1. 190/7 = 7) 190 2. 88/8 = 3. 762/7 = Decomposed these mixed numbers into improper fractions. To do this, multiply the denominator by the whole number, then add the numerator and put this number over the denominator. 4. 4 3/5 = 5. 5 1/9 = 6. 9 8/12 = Add these fractions together. Find a common denominator first. Two strategies to do this are: 1. Look to see if the smaller denominator number will go into the larger denominator number 2. Multiply the two denominators together Use 9 as the common denominator for this problem 7. 9 4/9 +2 2/3 Use 12 as the common denominator for this problem 8. 3 2/12 + 1 5/6 Use 32 for the common denominator for this problem 9. 6 3/8 + 1 1/4 Use 30 for the common denominator for this problem 10. 6 1/5 + 1 3/6 You find the common denominator for this problem 11. 4 7/8 +1 ΒΌ You find the common denominator for this problem 12. 6 3/8 +4 1/5

Extrinix:

First Portion: 1: How many times can 190 go into 7? 27 times, leaving 1 left so, \(27\dfrac{1}{7}\) 2: How many times can 88 go into 8? 11 times, so, \(11\) 3: How many times can 762 go into 7? 108 times, leaving 6 left, so \(108\dfrac{6}{7}\)

Extrinix:

Second Portion: \(\dfrac{whole~number \times denominator + numerator}{denominator}\) 4: \(4\times 5+3\) = \(\dfrac{23}{5}\) 5: \(5\times 9+1\) = \(\dfrac{46}{9}\) 6: \(9\times 12+8\) = \(\dfrac{116}{12}\)

Extrinix:

Third Portion: 7: Let's look for a similar denominator, well, how many times can 3 go into 9? 3 times, so multiply \(2\dfrac{2}{3}\) by 3, \(\dfrac{8}{3} \times 3\) = \(\dfrac{24}{9}\) = \(2\dfrac{6}{9}\) Now add these together, \(2\dfrac{6}{9}\) + \(9\dfrac{4}{9}\) = \(11\dfrac{10}{9}\) = \(12\dfrac{1}{9}\) 8: Look for a similar denominator, how many times can 6 go into 12? 2 times, so multiply \(1\dfrac{5}{6}\) by 2, \(\dfrac{11}{6} \times 2\) = \(\dfrac{22}{12}\) = \(1\dfrac{10}{12}\) Now add them together, \(1\dfrac{10}{12}\) + \(3\dfrac{2}{12}\) = \(4\dfrac{12}{12}\) = \(5\) 9: For this one, common denominator is 32, so multiply each fraction by the other fraction's opposite, \(\dfrac{51}{8} \times 4\) = \(\dfrac{204}{32}\) \(\dfrac{5}{4} \times 8\) = \(\dfrac{40}{32}\) Add together, \(\dfrac{204+40}{32}\) = \(\dfrac{244}{32}\) = \(7\dfrac{20}{32}\) 10: Multiply denominators, \(6\dfrac{1}{5} \times 6\) = \(\dfrac{186}{30}\) \(1\dfrac{3}{6} \times 5\) = \(\dfrac{45}{30}\) Add together, \(\dfrac{186+45}{30}\) = \(\dfrac{231}{30}\) = \(7\dfrac{21}{30}\) 11: How many times can 4 go into 8? 2 times, so multiply by 2, \(\dfrac{5}{4}\times 2\) = \(\dfrac{10}{8}\) = \(1\dfrac{2}{8}\) Add together, \(1\dfrac{2}{8}\) + \(4\dfrac{7}{8}\) = \(5\dfrac{9}{8}\) = \(6\dfrac{1}{8}\) 12: Multiply each fraction by other fractions denominator, \(6\dfrac{3}{8} \times 5\) = \(\dfrac{255}{40}\) \(4\dfrac{1}{5} \times 8\) = \(\dfrac{168}{40}\) Add together, \(\dfrac{255+168}{40}\) = \(\dfrac{423}{40}\) = \(10\dfrac{23}{40}\)

Extrinix:

If you need any more help on these feel free to dm me.

Extrinix:

Oh, or tag me in any posts if I'm on.